# JSJ decompositions of knot exteriors, Dehn surgery and the $L$-space conjecture

### v3

@@ 3.3. $LO$ and $NLS$ surgery on satellite knots and JSJ-graphs
Figure 1. $T_{1}$ and $T_{2}$ are two “swallow-follow” tori in the knot exterior of $K=K_{1}\#K_{2}$. Though $\partial X(K)$ is not drawn in the figure, it is obvious that $T_{1}$, $T_{2}$ and $\partial X(K)$ cobound a submanifold in $X(K)$. In this case, the submanifold is homeomorphic to the product of a pair of pants and $S^{1}$.
Figure 1 illustrates that the connected sum $K$ of two non-trivial knots contains two essential swallow-follow tori $T_{1},T_{2}$ which together with $\partial X(K)$ cobound a connected submanifold of $X(K)$. Note, moreover, that these tori have winding number $1$, so $K$ is not a cable knot. Hence as a special case of Theorem 3.10 we deduce that all rational surgeries on $K$ are $NLS$ and $LO$. This gives another proof of Krcatovich’s result that $L$-space knots are prime [38] and its $LO$ counterpart. In summary, we have the following corollary from Theorem 3.10.
@@
All rational surgeries on a composite knot are $LO$ and $NLS$. ∎
@@
We claim that $Y(r)$ is irreducible. Otherwise, the irreducibility of $K(r)$ implies that there exists a $2$-sphere in $Y(r)$ that bounds a $3$-ball $B$ in $K(r)$ but not in $Y(r)$. Then $B$ must contain some $T_{i}$, say $T_{1}$, and therefore $T_{1}$ is compressible in $K(r)$. Let $P_{1}$ denote $K$ viewed as a knot in $V_{1}$. Then $K(r)=X_{1}\cup_{T_{1}}P_{1}(r)$, where $P_{1}(r)=Y(r)\cup_{T_{2}}X_{2}$. Since $T_{1}$ is incompressible in $X_{1}$, it is compressible in $P_{1}(r)$ so one of the three possibilities listed in Theorem 2.9 applies to $P_{1}$ in $V_{1}$. Theorem 2.9(c) clearly doesn’t hold, and (b) doesn’t hold by our hypothesis that $r$ is not a cabling slope. Hence Theorem 2.9(a) holds and so $P_{1}$ is braided in $V_{1}$. But then by [2, Lemma 3.1], the torus $T_{2}$ bounds a solid torus in $V_{1}$, contradicting the fact that it bounds $X_{2}$.
@@
Figure 2. The figure illustrates the case that $Y$ is atoroidal and the green vertex $v$, representing the submanifold $Y$ is the root of the tree. In general, there could be other branches at the vertex $v$ in addition to the two branches depicted in the figure. The root of the tree is contained in one of these additional branches. The submanifold $Y$ is just $X(K)\setminus{\text{int}(X_{1})\cup\text{int}(X_{2})}$, corresponding to the union of all these additional branches.
@@
Suppose that $K$ is a satellite $L$-space knot. Then the JSJ graph of $K$ is a rooted interval.∎
@@
Suppose that $K$ is a satellite knot which admits an irreducible rational surgery that is not $LO$. Then the JSJ graph of $K$ is a rooted interval.∎
@@ 4. Co-oriented taut foliations and surgeries on satellite knots
Suppose that the exterior $X(K)$ of a knot $K$ contains disjoint essential tori $T_{1},T_{2}$ which together with $\partial X(K)$ cobound a connected submanifold of $X(K)$, and each of which bounds the exterior of a non-trivial fibred knot. Then $K(r)$ is $CTF$ unless $r$ is the cabling slope when $K$ is a cable knot.
@@
As explained in §3.3, the hypotheses of Theorem 4.2 hold if $K$ is a composite fibred knot. Hence we have the following corollary, which was previously proven by Delman and Roberts in [13] using different methods.
@@
(Delman-Roberts [13, Corollary 6.2]) All rational Dehn surgeries on a composite fibred knot are $CTF$. ∎
@@
In [8], we defined a slope $\alpha$ on the boundary of a knot manifold $M$ to be strongly $CTF$-detected if there exists a co-orientable taut foliation on $M$ that transversely intersects $\partial M$ in a linear foliation by simple closed curves of slope $\alpha$. If $\alpha$ is strongly $CTF$-detected, it is clear that it is $CTF$-detected and, moreover, that $M(\alpha)$ is $CTF$. It is unknown that whether $M(\alpha)$ being $CTF$ implies that $\alpha$ is strongly $CTF$-detected (cf. Remark 2.1).
@@
Let $K(r)=P(r)\cup_{T_{0}}X_{0}$ as before. Note that the longitudional slope $\lambda_{P(r)}$ of $P(r)$ is the slope $r/w^{2}$ on $\partial X_{0}$ by (2.2.6). If $T_{0}=\partial X_{0}$ is incompressible in $K(r)$, then $K(r)$ is $CTF$ by Lemma 2.6. Suppose, on the other hand, that $T_{0}$ is compressible in $K(r)$. Since we have assumed that $r$ is not a cabling slope, Theorem 2.13(1) and (2)(c) imply that $K(r)\cong K_{0}(r/w^{2})$ and therefore is $CTF$ by the hypothesis that $r/w^{2}$ is strongly $CTF$-detected. ∎
We note that the proof of Proposition 4.4 shows that if $\partial X_{0}$ is incompressible in $K(r)$, we need only assume that $r/w^{2}$ is $CTF$-detected in $X_{0}$ to conclude that $K(r)$ is $CTF$.
@@
Suppose that $K$ is a satellite knot with a persistently foliar companion whose associated pattern has winding number $w\geq 1$. Then $K(r)$ is $CTF$ for any $r\in\mathbb{Q}$ which is not the cabling slope if $K$ is a cable knot. ∎
@@
Lastly, we use Proposition 4.4 to deduce Proposition 4.7 below, which is based on a result of Roberts but also strengthens it. More precisely, in [50, Theorem 4.7], Roberts shows that if $K$ is a fibred knot whose monodromy has non-negative fractional Dehn twist coefficient, then any slope $r\in(-\infty,1)$ is strongly $CTF$-detected in $X(K)$ [^4] Though the statement of Theorem 4.7 in [50] requires the knot to be hyperbolic, the argument in [50] holds for all non-trivial fibred knots (cf. [13, Theorem 1.4]).. This, together with Proposition 4.4, immediately implies the following.
@@
Let $K$ be a satellite knot with a fibred companion whose associated winding number is $w$. Suppose that the fractional Dehn twist coefficent of the companion knot is non-negative. Then $K(r)$ is $CTF$ for each rational $r\in(-\infty,w^{2})$ unless $K$ is a cable knot and $r$ is the cabling slope. In particular this holds for fibred strongly quasipositive satellite knots.
@@
Let $K=P(K_{0})$ where $K_{0}$ is fibred. If $w=0$, then Theorem 4.1 shows that $K(r)$ is $CTF$ for any rational slope $r$ that is not a cabling slope. Now we assume that $w\geq 1$ and the fractional Dehn twist coefficent of the monodromy of $K_{0}$ is non-negative. By [50, Theorem 4.7], any slope in $(-\infty,1)$ is strongly $CTF$-detected. Then $K(r)$ is $CTF$ for any slope $r\in(-\infty,w^{2})$ that is not a cabling slope by Proposition 4.4.
Lastly, we explain why fibred strongly quasipositive satellite knots satisfy the hypotheses.
Let $K=P(K_{0})$ be a fibred strongly quasipositive knot. Then $K_{0}$ is fibred, and it is strongly quasipositive by the main result of [51]. Its monodromy is right-veering by [30, 33] and hence the fractional Dehn twist coefficient of its monodromy is non-negative. ∎
Remark 4.8.
Let $X_{0}$ be the exterior of the companion knot of the satellite knot $K$ in Proposition 4.7. The slope $1$ is $CTF$-detected in $X_{0}$ by Theorem 2.4. Then Lemma 2.6 implies that the conclusion of Proposition 4.7 holds for the slope $w^{2}$, and hence for all slopes $r\in(-\infty,w^{2}]$, if $\partial X_{0}$ is incompressible in $K(w^{2})$. This will be used in the proof of Theorem 4.11 below.
Since positive $L$-space knots (i.e. knots with a positive $L$-space surgery) are fibred and strongly quasipositive [45, 30], Proposition 4.7 applies to positive satellite $L$-space knots. In fact, Lemma 4.10 below shows that in that case, the possibility that $r$ is a cabling slope does not arise. Hence we have the following corollary.
@@
If $K$ is a positive satellite $L$-space knot with pattern of winding number $w$ then $K(r)$ is $CTF$ for each rational $r\in(-\infty,w^{2})$. ∎
@@
Let $K=P(K_{0})$ be a positive satellite $L$-space knot where $P$ is a cabled pattern with winding number $w$. Then the cabling slope of $K$ is greater than or equal to $w^{2}$.
@@
We first consider the case $P=C_{m,n}$. By [32], the assumption that $K$ is a positive $L$-space knot implies that $n/m\geq 2g(K_{0})-1\geq 1$. So $n\geq m$. Then the cabling slope of $K=mn\geq m^{2}=w^{2}$.
Now suppose $P=C_{m,n}(P_{0})$ where $P_{0}$ is a pattern. Let $K_{1}=P_{0}(K_{0})$; so $K=C_{m,n}(K_{1})$. Let $w_{0}=w(P_{0})$ and $g_{0}=g(K_{0})$. Then by [53],
$g(K_{1})=g(P_{0}(U))+w_{0}g_{0}$
Since $K_{1}$ is also a satellite $L$-space knot by [27, Theorem 1.15] (cf. [32, Proposition 3.3]). By [1, Theorem 7.5(1)] [^5] Theorem 7.5 in [1] is based on the assumption that [1, Conjecture 7.1], originally posed by J. Rasmussen and S. Rasmussen in [48], holds. In this article we work with Heegaard Floer homology with $\mathbb{Z}_{2}$-coefficents, and the $\mathbb{Z}_{2}$-version of [1, Conjecture 7.1] has been established by Hanselman, Rasmussen and Watson in [27, Theorem 1.14]., $2g(P_{0}(U))-1+w_{0}>w_{0}(w_{0}-1)(2g_{0}-1)$ and therefore
$1\displaystyle 2g(K_{1})-$ 1 $=$ $2g(P_{0}(U))-1+2w_{0}g_{0}$
$>$ $w_{0}(w_{0}-1)(2g_{0}-1)-w_{0}+2w_{0}g_{0}$
$=$ $w_{0}^{2}(2g_{0}-1)$
$\geq$ $w_{0}^{2}$
Hence as $n/m\geq 2g(K_{1})-1>w_{0}^{2}$ by [32], $mn>m^{2}w_{0}^{2}=w^{2}$. ∎
@@
Let $K=P(K_{0})$ be a positive satellite $L$-space knot. Then $K(r)$ is $CTF$ if $r\leq 9$ unless $K(r)$ is an $L$-space. The latter happens exactly when $K=C_{2,n}(T(2,3))$ and either
  1. (1)
    $n=3,r\in[5,9]$ or
  2. (2)
    $n=5,r\in[7,9]$ or
  3. (3)
    $n=7,r=9$.
@@
Let $w$ denote the winding number of $P$. When $w>3$, the result follows from Corollary 4.9. So we can assume that $w\in\{2,3\}$ by Corollary 3.4.
If $w=3$, Corollary 4.9 implies that $K(r)$ is $CTF$ for $r<9$. The same conclusion holds when $r=9$ as long as $T$ is incompressible in $K(9)$ by Remark 4.8. But if $T$ compresses, Theorem 2.13 implies $P=C_{3,n}$ for an integer $n$ coprime with $3$ such that $|9-3n|=\Delta(3n/1,9/1)\leq 1$, which is impossible.
The final case to consider is when $w=2$. Since $K$ is a positive $L$-space knot, the companion $K_{0}$ is a positive $L$-space knot by [27, Theorem 1.15], and $P$ is braided in its solid torus by [1, Theorem 7.4] (also see footnote $5$). So $P=C_{2,n}$ for some odd $n$. We consider three cases: $g(K_{0})=1$, $g(K_{0})=2$ and $g(K_{0})>2$.
@@
Then $K_{0}=T(2,3)$ by [24]. Since $K$ is a positive $L$-space knot, [32] implies that $n/2\geq 2g(K_{0})-1=1$, so as $n$ is odd, $n\geq 3$. Also, $C_{2,n}(U)=T(2,n)$, and therefore $g(K)=g(C_{2,n}(U))+2g(K_{0})=(n-1)/2+2=(n+3)/2$ by [53]. If $r\neq 2n$, then $K(r)$ is either Seifert fibred or a graph manifold, therefore it is $CTF$ if and only if it is $NLS$ ([41, 7, 3, 26]), and the latter happens if and only if $r<2g(K)-1=n+2$ by [46, 47]. For $n+2\leq r\leq 9$, we obtain the exceptions listed in the theorem. When $r=2n\leq 9$, we have $n=3$ and $r=2n=6$. So $K(6)=L(2,3)\#T(2,3)(3/2)$ is an $L$-space and is not $CTF$.
@@
Then $K_{0}=T(2,5)$ by [14]. By [32], $n/2\geq 2g(K_{0})-1=3$, so $n\geq 7$ and the cabling slope $2n\geq 14$. Also, $g(K)=(n-1)/2+2\cdot 2=(n+7)/2$, and hence, as in the case above, $K(r)$ is $CTF$ if and only if $r<2g(K)-1=n+6$ by [46, 47, 41, 7, 3, 26]. Therefore $K(r)$ is $CTF$ for $r<13$.
@@
In this case, we will show that $K(r)$ is $CTF$ for any $r<13$. By [32], $n/2\geq 2\cdot 3-1=5$, so $n\geq 11$ and the cabling slope is $2n\geq 22$. We have $K(r)=C_{2,n}(r)\cup_{T}X_{0}$, where $X_{0}$ is the exterior of $K_{0}$.
@@
Therefore there is a slope $s$ on $T$ that is $CTF$-detected in both $C_{2,n}(r)$ and $X_{0}$. Hence $K(r)$ is $CTF$ for $r<13$.
@@
If $K$ is a positive satellite $L$-space knot then for any $r\leq 9$, $K(r)$ is $CTF$ if and only if it is $NLS$. ∎

### v4

@@ 3.3. $LO$ and $NLS$ surgery on satellite knots and JSJ-graphs
Figure 1. $T_{1}$ and $T_{2}$ are two “swallow-follow” tori in the exterior of $K=K_{1}\#K_{2}$. Though $\partial X(K)$ is not drawn in the figure, it is obvious that $T_{1}$, $T_{2}$ and $\partial X(K)$ cobound a submanifold in $X(K)$. In this case, the submanifold is homeomorphic to the product of a pair of pants and $S^{1}$.
Figure 1 illustrates that the exterior $X(K)$ of the connected sum $K$ of two non-trivial knots contains two essential swallow-follow tori $T_{1},T_{2}$ which together with $\partial X(K)$ cobound a connected submanifold of $X(K)$. Note, moreover, that these tori have winding number $1$, so $K$ is not a cable knot. Hence as a special case of Theorem 3.10 we deduce that all rational surgeries on $K$ are $NLS$ and $LO$. This gives another proof of Krcatovich’s result that $L$-space knots are prime [38] and its $LO$ counterpart. In summary, we have the following corollary from Theorem 3.10.
@@
All rational surgeries on a composite knot are $LO$ and $NLS$.
@@
We claim that $Y(r)$ is irreducible. Otherwise, the irreducibility of $K(r)$ implies that there exists a $2$-sphere in $Y(r)$ that bounds a $3$-ball $B$ in $K(r)$ but not in $Y(r)$. Then $B$ must contain some $T_{i}$, say $T_{1}$, and therefore $T_{1}$ is compressible in $K(r)$. Let $P_{1}$ denote $K$ viewed as a knot in $V_{1}$. Then $K(r)=X_{1}\cup_{T_{1}}P_{1}(r)$, where $P_{1}(r)=Y(r)\cup_{T_{2}}X_{2}$. Since $T_{1}$ is incompressible in $X_{1}$, it is compressible in $P_{1}(r)$ so either case (a) or case (b) of Theorem 2.9 applies to $P_{1}$ in $V_{1}$. Case (b) doesn’t hold by our hypothesis that $r$ is not a cabling slope. Hence Theorem 2.9(a) holds and so $P_{1}$ is braided in $V_{1}$. But then by [2, Lemma 3.1], the torus $T_{2}$ bounds a solid torus in $V_{1}$, contradicting the fact that it bounds $X_{2}$.
@@
Figure 2. The figure illustrates the case that $Y$ is atoroidal and the green vertex $v$, representing the submanifold $Y$, is the root of the tree. In general, there could be other branches at the vertex $v$ in addition to the two branches depicted in the figure. The root of the tree is contained in one of these additional branches. The submanifold $Y$ is just $X(K)\setminus{\text{int}(X_{1})\cup\text{int}(X_{2})}$, corresponding to the union of all these additional branches.
@@
Suppose that $K$ is a satellite $L$-space knot. Then the JSJ graph of $K$ is a rooted interval.
@@
Suppose that $K$ is a satellite knot which admits an irreducible rational surgery that is not $LO$. Then the JSJ graph of $K$ is a rooted interval.
@@ 4. Co-oriented taut foliations and surgeries on satellite knots
Suppose that the exterior $X(K)$ of a knot $K$ contains disjoint essential tori $T_{1},T_{2}$ which together with $\partial X(K)$ cobound a connected submanifold of $X(K)$, and each of which bounds the exterior of a fibred knot. Then $K(r)$ is $CTF$ unless $K$ is a cable knot and $r$ is the cabling slope.
@@
An immediate corollary of Theorem 4.2 is the following $CTF$ analogue of Theorem 3.13.
Theorem 4.3.
Suppose that $K$ is a fibered satellite knot which admits an irreducible rational surgery which is not $CTF$. Then the JSJ graph of $K$ is a rooted interval.
It is interesting to note that since $CTF$ implies $NLS$, the only additional facts needed to deduce Theorem 3.12 from Theorem 4.3 are that an $L$-space knot is fibered and has at least one irreducible $L$-space rational surgery.
Theorem 4.3 has the following corollary, which was previously proven by Delman and Roberts in [13] using different methods.
@@
(Delman-Roberts [13, Corollary 6.2]) All rational Dehn surgeries on a composite fibred knot are $CTF$.
@@
In [8], we defined a slope $\alpha$ on the boundary of a knot manifold $M$ to be strongly $CTF$-detected if there exists a co-orientable taut foliation on $M$ that transversely intersects $\partial M$ in a linear foliation by simple closed curves of slope $\alpha$. If $\alpha$ is strongly $CTF$-detected, it is clear that it is $CTF$-detected and, moreover, that $M(\alpha)$ is $CTF$. It is unknown whether $M(\alpha)$ being $CTF$ implies that $\alpha$ is strongly $CTF$-detected (cf. Remark 2.1).
@@
Let $K(r)=P(r)\cup_{T_{0}}X_{0}$ as before. Note that the longitudinal slope $\lambda_{P(r)}$ of $P(r)$ is the slope $r/w^{2}$ on $\partial X_{0}$ by (2.2.6). If $T_{0}=\partial X_{0}$ is incompressible in $K(r)$, then $K(r)$ is $CTF$ by Lemma 2.6. Suppose, on the other hand, that $T_{0}$ is compressible in $K(r)$. Since we have assumed that $r$ is not a cabling slope, Theorem 2.13(1) and (2)(c) imply that $K(r)\cong K_{0}(r/w^{2})$ and therefore is $CTF$ by the hypothesis that $r/w^{2}$ is strongly $CTF$-detected. ∎
Remark 4.6.
The proof of Proposition 4.5 shows that if $\partial X_{0}$ is incompressible in $K(r)$, we need only assume that $r/w^{2}$ is $CTF$-detected in $X_{0}$ to conclude that $K(r)$ is $CTF$.
@@
Suppose that $K$ is a satellite knot with a persistently foliar companion whose associated pattern has winding number $w\geq 1$. Then $K(r)$ is $CTF$ for any $r\in\mathbb{Q}$ which is not the cabling slope if $K$ is a cable knot.
@@
Lastly, we use Proposition 4.5 to deduce Proposition 4.9 below, which is based on a result of Roberts but also strengthens it. More precisely, in [50, Theorem 4.7], Roberts shows that if $K$ is a fibred knot whose monodromy has positive (resp. negative, resp. zero) fractional Dehn twist coefficient, then any slope $r\in(-\infty,1)$ (resp. $(-1,\infty)$, resp. (,)($\infty$,$\infty$)( - ∞ , ∞ )) is strongly $CTF$-detected in $X(K)$ [^5] Though the statement of Theorem 4.7 in [50] requires the knot to be hyperbolic, the argument in [50] holds for all non-trivial fibred knots (cf. [13, Theorem 1.4]).. This, together with Proposition 4.5, implies the following.
@@
Let $K=P(K_{0})$ where $K_{0}$ is fibered and $P$ has winding number $w$. Let $c$ be the fractional Dehn twist coefficient of the monodromy of $K_{0}$. Suppose that $r$ is not the cabling slope if $K$ is a cable knot. Then $K(r)$ is $CTF$ if either
  1. (1)
    $c>0$ and $r\in(-\infty,w^{2}]$, or
  2. (2)
    $c<0$ and $r\in[-w^{2},\infty)$, or
  3. (3)
    $c=0$ and $r\in(-\infty,\infty)$.
@@
If $w=0$, then Theorem 4.1 shows that $K(r)$ is $CTF$ for any rational slope $r$ that is not a cabling slope. So we assume that $w\geq 1$.
If $c>0$ then by [50, Theorem 4.7], any slope in $(-\infty,1)$ is strongly $CTF$-detected in $X_{0}$. Then $K(r)$ is $CTF$ for any slope $r\in(-\infty,w^{2})$ that is not a cabling slope by Proposition 4.5.
To deal with the case $r=w^{2}$ we use the facts that $\partial X_{0}$ is incompressible in $K(w^{2})$ by Lemma 2.14, and that the slope 1 is $CTF$-detected in $X_{0}$ by Theorem 2.4. Then $K(w^{2})$ is $CTF$ by Remark 4.6.
This proves (1). (2) is completely analogous, and (3) follows in the same way using the fact that if $c=0$ then all rational slopes on $\partial X_{0}$ are strongly $CTF$-detected in $X_{0}$ [50, Theorem 4.7]. ∎
We note that fibred strongly quasipositive satellite knots satisfy the hypothesis (1). For, suppose $K=P(K_{0})$ is fibred and strongly quasipositive. Then $K_{0}$ is fibred, and it is strongly quasipositive by the main result of [51]. Its monodromy is right-veering by [30, 33] and hence the fractional Dehn twist coefficient of its monodromy is positive.
It is known that positive $L$-space knots (i.e. knots with a positive $L$-space surgery) are fibred and strongly quasipositive [45, 30]. We say that a satellite knot is a positive satellite $L$-space knot if it is also a positive $L$-space knot. Then Proposition 4.9(1) applies to positive satellite $L$-space knots. In fact, Lemmas 4.11 and 4.12 below show that in that case, the possibility that $r$ is a cabling slope does not arise. Hence we have the following corollary, which by Lemma 4.12 is predicted by the $L$-space Conjecture.
@@
If $K$ is a positive satellite $L$-space knot with pattern of winding number $w$ then $K(r)$ is $CTF$ for each rational $r\in(-\infty,w^{2}]$.
@@
Let $K$ be a positive satellite $L$-space cable knot. Then surgery on $K$ along the cabling slope is an $L$-space.
@@
We have $K=C_{m,n}(K_{0})$. Let $g_{0}=g(K_{0})$. By [32], the assumption that $K$ is a positive $L$-space knot implies that $n/m>2g_{0}-1\geq 1$, so $m+n>2mg_{0}$.
By [53],
$g(K)=g(C_{m,n}(U))+mg_{0}$
Since $C_{m,n}(U)$ is the torus knot $T(m,n)$ we get
$1\displaystyle 2g(K)-$ 1 $=(m-1)(n-1)+2mg_{0}-1$
$=mn-(m+n-2mg_{0})<mn$
Lemma 4.12.
Let $K$ be a positive satellite $L$-space knot with pattern of winding number $w$. Then $w^{2}<2g(K)-1$.
Proof.
We have $K=P(K_{0})$, where $P$ has winding number $w$. Let $g_{0}=g(K_{0})$.
By [53],
$g(K)=g(P(U))+wg_{0}$
Hence
$2g(K)-1=2g(P(U))-1+2wg_{0}$
By [1, Theorem 7.3(i)],
$1\displaystyle 2g(P(U))-$ 1 $>w(w-1)(2g_{0}-1)-w$
$=(w-1)2wg_{0}-w^{2}$
Hence $2g(K)-1>w^{2}(2g_{0}-1)\geq w^{2}$. ∎
@@
Let $K=P(K_{0})$ be a positive satellite $L$-space knot. Then $K(r)$ is $CTF$ if $r\in(-\infty,9]$ unless $K(r)$ is an $L$-space. The latter happens exactly when $K=C_{2,n}(T(2,3))$ with $n=3,5,$ or $7$ and $r\in[n+2,9]$.
@@
Let $w$ denote the winding number of $P$. When $w\geq 3$, the result follows from Corollary 4.10. So we can assume that $w=2$ by Corollary 3.4.
Since $K$ is a positive $L$-space knot, the companion $K_{0}$ is a positive $L$-space knot by [27, Theorem 1.15], and $P$ is braided in its solid torus by [1, Theorem 1.17] (also see footnote $6$). So $P=C_{2,n}$ for some odd $n$. We consider three cases: $g(K_{0})=1$, $g(K_{0})=2$ and $g(K_{0})>2$.
@@
Then $K_{0}=T(2,3)$ by [24]. Since $K$ is a positive $L$-space knot, [32] implies that $n/2>2g(K_{0})-1=1$, so $n\geq 3$. Also, $C_{2,n}(U)=T(2,n)$, and therefore $g(K)=g(C_{2,n}(U))+2g(K_{0})=(n-1)/2+2=(n+3)/2$ by [53]. If $r\neq 2n$, then $K(r)$ is either Seifert fibred or a graph manifold, therefore it is $CTF$ if and only if it is $NLS$ ([41, 7, 3, 26]), and the latter happens if and only if $r<2g(K)-1=n+2$ by [46, 47]. For $n+2\leq r\leq 9$, we obtain the exceptions listed in the theorem. When $r=2n\leq 9$, we have $n=3$ and $r=2n=6$. So $K(6)=L(2,3)\#T(2,3)(3/2)$ is an $L$-space and is not $CTF$.
@@
Then $K_{0}=T(2,5)$ by [14]. By [32], $n/2>2g(K_{0})-1=3$, so $n\geq 7$ and the cabling slope $2n\geq 14$. Also, $g(K)=(n-1)/2+2\cdot 2=(n+7)/2$, and hence, as in the case above, $K(r)$ is $CTF$ if and only if $r<2g(K)-1=n+6$ by [46, 47, 41, 7, 3, 26]. Therefore $K(r)$ is $CTF$ for $r<13$.
@@
In this case, we will show that $K(r)$ is $CTF$ for any $r<13$. By [32], $n/2>2\cdot 3-1=5$, so $n\geq 11$ and the cabling slope is $2n\geq 22$. We have $K(r)=C_{2,n}(r)\cup_{T}X_{0}$, where $X_{0}$ is the exterior of $K_{0}$.
@@
Therefore there is a slope $s$ on $T$ that is $CTF$-detected in both $C_{2,n}(r)$ and $X_{0}$. Hence $K(r)$ is $CTF$ for $r<13$. ∎
@@
If $K$ is a positive satellite $L$-space knot then for any $r\in(-\infty,9]$, $K(r)$ is $CTF$ if and only if it is $NLS$.